# The Effect of Tornadic Supercell Thunderstorms on the Atmospheric Muon Flux

### v3

@@ Abstract
Tornadoes are severe weather phenomena characterized by a violently rotating column of air connecting the ground to a parent storm. Within the United States, hundreds of tornadoes occur every year. Despite this, the dynamics of tornado formation and propagation are not particularly well understood, in part due to the challenge of instrumentation: many existing instruments for measuring atmospheric properties are in-situ detectors, making deployment in or near an active or developing tornado difficult. Here, we combine local atmospheric and cosmic ray air shower simulation to explore the potential for remote measurement of the pressure field within tornado-producing supercell thunderstorms by examining directional variations of the atmospheric muon flux.
@@ I.2 Atmospheric Muons
where E is the energy of the muon, $x$ is the distance traveled, $\rho(x)$ is the density of the traversed matter and $a(x,E)$, $b(x,E)$ are the ionisation and radiative energy losses. The quantities $\rho$, $a$, and $b$ are all dependent on the properties of the matter traversed. For atmospheric muons, $\rho$ is directly proportional to the atmospheric pressure [7].
@@
The atmospheric muon flux is known to be proportional to the local atmospheric pressure. This has been demonstrated to be true on large scales: daily and weekly variations in the atmospheric muon flux that are correlated with atmospheric pressure have been observed by various muon detectors [8, 9]. On smaller scales, local variations in atmospheric pressure have been observed in data from muon telescopes pointed at typhoons off the coast of Japan [10], as well as non-tornadic thunderstorms [11, 12].
@@ II.1 Conceptual Overview
Since the atmospheric muon flux is correlated with the density of matter along the muon path, and the density of air in the atmosphere is correlated with the atmospheric air pressure, this implies that variations in atmospheric air pressure can be observed as variations in the atmospheric muon flux. A hypothetical muon detector could detect variations in air pressure as a function of altitude by measuring the muon flux at different zenith angles. As mentioned above, this has been done to measure the low pressure region of typhoons, where a region of increased muon flux was found to be correlated with the low-pressure regions of the target typhoon [10]. Given the success of this method in the context of typhoon measurements, we consider extending this technique to make a similar measurement of the pressure field within supercell thunderstorms.
@@
  • Given an estimate of the pressure perturbations within a supercell thunderstorm, how large is the corresponding effect on the atmospheric muon flux?
@@
Figure 1: Left: A cartoon of the setup considered in this paper (not to scale). Cosmic rays interact in the upper atmosphere, and secondary muons propagate through a supercell thunderstorm. A muon detector is placed on the opposite side of the storm, and the effect of the thunderstorm pressure field on the flux of atmospheric muons is calculated. Right: A top-down view of the simulated arrangement, showing the location of the muon detector as a red “x”, as well as the simulated storm radar reflectivity at an altitude of 1.54 km. The tornado is located in the hook echo near $(x,y)=(0,10)$
@@ II.2 Supercell Thunderstorm Simulation
Cloud Model 1 (CM1), a three-dimensional nonhydrostatic fully compressible cloud model, was used to simulate the tornadic supercell thunderstorm [15]. Among the prognostic variables CM1 solves for include wind, potential temperature, pressure, turbulent kinetic energy, and water substance including water vapor, cloud water, cloud ice, rain, snow, graupel and hail. For the purpose of this study, only the pressure field was analyzed. The initial environmental conditions for the simulation were identical to [16] and [17], namely, the state of the atmosphere prior to the observed 24 May 2011, El Reno, Oklahoma supercell that produced a long-track violent tornado. Physics parameterization options for the simulation include three-moment microphysics [18] and a subgrid turbulent kinetic energy closure scheme [19]. The simulation utilized an isotropic cubic mesh with a grid spacing of 75 m, and was triggered using an updraft nudging technique [20]. The simulation morphology is similar to [16], with the storm initially splitting with a dominant right-moving supercell (analyzed herein) producing a vigorous low-level rotating updraft (mesocyclone) prior to the formation of a violent tornado. The pressure field of the storm is analyzed at six model times, detailed in table 1. A tornado forms at $t=5580$ seconds, producing a region of low density (figure 3), and a region of high vertical vorticity (seen in figure 4).
@@ II.3 Muon Simulation
MCEq [^1] https://github.com/mceq-project/MCEq/tree/master/MCEq is used to numerically solve the atmospheric particle cascade equations describing the muon flux as the shower propagates through the atmosphere. The primary cosmic ray flux is simulated according to the model described in [21]. Primary cosmic ray interactions are then simulated using the SIBYLL 2.3c interaction model [22]. The resultant muon flux is then propagated through an atmospheric model described in the previous section.
@@
where $\Delta\theta$ and $\Delta\phi$ correspond to the size of the angular bins in azimuth and elevation (2 degrees). The bounds of the energy integral, $E_{min}$ and $E_{max}$ describe the minimum and maximum muon energies that were simulated. For this paper, muon fluxes were calculated between $E_{min}=0\text{ TeV}$ and $E_{max}=10^{8}\text{ TeV}$, however the vast majority ($>99\%$) of the atmospheric muon flux reching the ground exists between 0 and 1 TeV. We choose $\Delta t$ = 1 hour, examining the storm at set times using simulated muon data from the preceding hour. This ensures a large quantity of muon detections, while also being a short enough time interval that the evolution of the storm can be studied. To give a sense of scale: for a 1000 square meter muon detector, the expected rate of atmospheric muons is approximately 28,400 events per square degree angular bin per hour at $\theta=40^{\circ}$. This rate varies naturally as a function of elevation angle, with a lower flux near the horizon and a higher flux from very vertical elevation angles [27].
@@ III Results
Figure 2 shows the idealized expectation of the effect of the storm density field on the atmospheric muon flux along various lines of sight for a 1000 m2 muon detector, over the course of 1 hour preceding the time listed for the map. Regions of increased muon flux, identifiable as groups of pixels with higher values of $L$, correspond to low-density regions of the storm. Similarly, regions of suppressed muon flux correspond to higher-density regions. Note, however, that the muon flux is correlated with the average of the density field along a particular line of sight, meaning information about the density field as a function of depth is not easily accessible from the two dimensional images shown. Depth-dependent information about the density field could conceivably be obtained by multiple detectors with different views of the same storm, or by examining the energy distribution of detected muons, however development of this idea is left for future study.
@@
For comparison, plots showing the density field integrated along lines of sight leading to the muon detector are shown in figure 3. The large region of low density located near $\phi=100$ degrees corresponds to the storm’s mesocyclone. While the mesocyclone is present during all stages of the storm’s lifecycle, the intensification of this region between $t=5580$ and $t=10200$ seconds corresponds to the presence of a tornado. This intensification is reflected in the muon flux, as can be seen in plots of the idealized muon flux in figure 2.
Figure 3: The density field integrated along line of sight leading to the muon detector (here, $d\ell$ is the line element along the line of sight), at various times during the simulated storm evolution. The tornado forms at $t=5580$ seconds, and is located within the mesocyclone producing the density perturbation near $\phi=100$ degrees.
@@
The plots show in figure 2 show the idealized expectation of the effect of the simulated supercell thunderstorm on the atmospheric muon flux, neglecting statistical variations in the atmospheric muon flux. To improve upon this, Poisson variations are added to $N_{obs}$ and $L$ is recalculated using the new values. Examples of maps of $L$ containing these variations can be seen in figure 5. Near the horizon, the muon flux is lower and the data is dominated by Poisson variations. At high elevations ($\theta>30$ degrees) the structure of the high and low pressure regions of the storm begin to become visible as clusters of pixels with lower/elevated $L$ values. The intensification of the low pressure core temporally coincident with tornado formation is still visible, even when accounting for statistical variations in the observed muon flux over a 1 hour time window.
The effect of the tornado can also be seen in the total summed value of the maps shown in figure 5. Low pressure perturbations will introduce an excess of muon events seen by the detector, resulting in lower per-pixel $\log(L)$ values on average. Figure 6 shows the value of the summed muon images as a function of time during the supercell thunderstorm’s life span. Tornado formation introduces a drop in the spatially integrated muon flux beyond the effect seen from the mesocyclone alone.
@@ IV.1 Muon Detector Size
Smaller detectors may still be able to measure atmospheric pressure effects on the muon flux. Since variations in the atmospheric muon flux due to supercell thunderstorms are on the order of 1%, muon detectors seeking to measure this effect need to be large enough to collect O($10^{6}$) atmospheric muon counts over the angular area of interest during the lifetime of the storm (typically 10s to 100s of minutes). Measurements taken with smaller detectors will have larger relative statistical fluctuations, making it more difficult to identify the underlying structure of the pressure field. On the other hand, smaller detectors also have the potential to be significantly more portable. While a 1000 m2 would likely need to be a permanent, stationary installation, a 100 m2 detector could conceivably be moved to a location with projected severe weather several days beforehand, and a 50 m2 detector is small enough to fit on roughly 5 trucks, allowing for the detector to be moved to the location of a severe weather event as it is developing. Examples of muon maps produced by a 100 m2 detector can be seen in figure 7. Though statistical variations conceal much of the structure seen in figure 5, the aggregate effect of the tornado on the underlying pressure field can still be observed in the sum of the individual pixel values, albeit to a much smaller degree. This is shown in figure 8.
Figure 7: Maps of the muon flux, similar to those shown in figure 5, but as seen by a 100 m2 muon detector. The effect of the low-pressure core of the tornado is significantly smaller relative to statistical fluctuations in comparison to the 1000 m2 case, though is still faintly visible.
@@ IV.2 Range and Resolution
Though individual pressure measurements using this technique may not be as precise in measuring the pressure at a particular location when compared to more traditional in-situ methods, there is a distinct advantage in being able to remotely measure the pressure field over large distance scales. Measurements of the atmospheric muon flux can reveal unique information of the variation of the pressure field as a function of height that is inaccessible to ground-based pressure sensors. Additionally, the lack of a need for in-situ instrumentation makes this measurement more logistically straightforward than current methods, as it does not require detectors to be within the direct path of an active tornado.
The range of this technique depends on the desired spatial resolution of the pressure perturbation field. The atmospheric muon flux naturally varies with elevation angle, with a higher proportion of muon events incident at large elevation angles [27]. For a fixed detector size, a more detailed scan of the pressure field requires a larger number of muon counts, which can be obtained by using the muon flux a higher (more vertical) elevation angles. However, using data at higher elevation angles reduces the maximum possible horizontal distance between the detector and the storm.
@@
where $d$ is the distance scale over which pressure variations are measured, $\theta$ is the elevation angle of the target region, and $\Delta\psi$ is obtained for a particular elevation angle from the curves shown in figure 9. Due to the competing effects of the increased resolution at low elevation angles, and the decreased horizontal distance when examining high elevation angles, the horizontal range of this technique peaks near an elevation angle of 45 degrees. This can be seen for various detector sizes and desired distance resolutions in figure 10. Note that here “horizontal range” is the distance along the surface of the earth, accounting for the earth’s curvature.
With a 1000 m2 muon detector, a 5 kilometer resolution measurement of the pressure field could be obtained at a distance of up to 35 kilometers, though this is restricted to along the $\theta=40^{\circ}$ line of sight. If a larger, permanent muon detector were placed in a tornado-prone location, this would translate to being able to measure the pressure field of $O(1)$ tornadoes every year [30]. Smaller detectors would have a reduced, but still significant range, capable of measuring pressure variations over 5 kilometer length scales at a range of 10 kilometers. Such detectors would have the important advantage of portability, with the potential to be deployed or moved near regions expected to experience severe weather in the near future.
@@
Figure 10: Projected horizontal ranges of using this technique to measure pressure field variations along the listed length scales ($d$). Large elevation angles correspond to a higher muon rate allowing for more precise measurements, but also geometrically correspond to a smaller horizontal distance to a particular altitude layer. These two competing effects are roughly balanced near an elevation angle of 45 degrees, where the horizontal range of this technique is maximized.
@@ IV.4 Other Weather Phenomena
In the more general case, atmospheric phenomena can be characterized by their physical size and the scale of their pressure perturbations. By simulating various atmospheric phenomena described by these two variables, we can estimate what types of atmospheric phenomena are candidates for remote pressure measurements using atmospheric muons.
A simplified simulation of the pressure field is used to achieve this. Different atmospheric phenomena are simulated as 10-kilometer tall cylinders of negative pressure perturbation, with variable horizontal radius. The radius and average pressure perturbation are varied, and for each case the muon flux residual integrated over the full field of view is recorded. This can be used to estimate how large of an effect various atmospheric phenomena might have on the muon flux. These curves, as well as highlighted regions corresponding to different atmospheric phenomena, can be seen in figure 12. These results seem consistent with previous studies [10, 12], which have observed perturbations in the muon flux in the vicinity of typhoons (hurricanes) and non-tornadic thunderstorms.
Figure 12: Various severe weather phenomena plotted in the space of size (x-axis), and magnitude of associated pressure drop (y-axis) [31, 32, 14, 33, 34]. Highlighted regions for tornadoes, thunderstorms, and hurricanes are approximate, as these values have not been exhaustively cataloged for the majority of storms. For the purposes of this plot, supercell thunderstorms are estimated to have pressure deviations comparable to that of tornadoes, though this has yet to be measured. Curves corresponding to regions of a 1% and 5% effect on the atmospheric muon flux are shown, where points above these curves are expected to produce a larger deviation in the muon flux.

### v5

@@ Abstract
Tornadoes are severe weather phenomena characterized by a violently rotating column of air connecting the ground to a parent storm. Within the United States, hundreds of tornadoes occur every year. Despite this, the dynamics of tornado formation and propagation are not particularly well understood, in part due to the challenge of instrumentation: many existing instruments for measuring atmospheric properties are in-situ detectors, making deployment in or near an active or developing tornado difficult. Here, we combine local atmospheric and cosmic ray air shower simulation to explore the potential for remote measurement of the density field within tornado-producing supercell thunderstorms by examining directional variations of the atmospheric muon flux.
@@ I.2 Atmospheric Muons
where E is the energy of the muon, $x$ is the distance traveled, $\rho(x)$ is the density of the traversed matter and $a(x,E)$, $b(x,E)$ are the ionisation and radiative energy losses. The quantities $\rho$, $a$, and $b$ are all dependent on the properties of the matter traversed. For atmospheric muons, $\rho$ is directly proportional to the atmospheric pressure [7], suggesting that muographic measurements of air density could be used to infer the atmospheric pressure in a region.
@@
The atmospheric muon flux is known to be proportional to the local atmospheric pressure (and thus, atmospheric density as well). This has been demonstrated to be true on large scales: daily and weekly variations in the atmospheric muon flux that are correlated with atmospheric pressure have been observed by various muon detectors [8, 9]. On smaller scales, local variations in atmospheric pressure have been observed in data from muon telescopes pointed at typhoons off the coast of Japan [10], as well as non-tornadic thunderstorms [11, 12].
@@ II.1 Conceptual Overview
Since the atmospheric muon flux is correlated with the density of matter along the muon path, and the density of air in the atmosphere is correlated with the atmospheric air pressure, this implies that variations in atmospheric air pressure can be observed as variations in the atmospheric muon flux. A hypothetical muon detector could detect variations in air pressure as a function of altitude by measuring the muon flux at different zenith angles. As mentioned above, this has been done to measure the low pressure region of typhoons, where a region of increased muon flux was found to be correlated with the low-pressure regions of the target typhoon [10]. Given the success of this method in the context of typhoon measurements, we consider extending this technique to make a similar measurement of the density field within supercell thunderstorms.
@@
  • Given an estimate of the density perturbations within a supercell thunderstorm, how large is the corresponding effect on the atmospheric muon flux?
@@
Figure 1: Left: A cartoon of the setup considered in this paper (not to scale). Cosmic rays interact in the upper atmosphere, and secondary muons propagate through a supercell thunderstorm. A muon detector is placed on the opposite side of the storm, and the effect of the thunderstorm density field on the flux of atmospheric muons is calculated. Right: A top-down view of the simulated arrangement, showing the location of the muon detector as a red “x”, as well as the simulated storm radar reflectivity at an altitude of 1.54 km. The tornado is located in the hook echo near $(x,y)=(0,10)$
@@ II.2 Supercell Thunderstorm Simulation
Cloud Model 1 (CM1), a three-dimensional nonhydrostatic fully compressible cloud model, was used to simulate the tornadic supercell thunderstorm [15]. Among the prognostic variables CM1 solves for include wind, potential temperature, pressure, turbulent kinetic energy, dry air density, total density and water substance including water vapor, cloud water, cloud ice, rain, snow, graupel and hail. For the purpose of this study, only the total density field was analyzed. The initial environmental conditions for the simulation were identical to [16] and [17], namely, the state of the atmosphere prior to the observed 24 May 2011, El Reno, Oklahoma supercell that produced a long-track violent tornado. Physics parameterization options for the simulation include three-moment microphysics [18] and a subgrid turbulent kinetic energy closure scheme [19]. The simulation utilized an isotropic cubic mesh with a grid spacing of 75 m, and was triggered using an updraft nudging technique [20]. The simulation morphology is similar to [16], with the storm initially splitting with a dominant right-moving supercell (analyzed herein) producing a vigorous low-level rotating updraft (mesocyclone) prior to the formation of a violent tornado. The density field of the storm is analyzed at six model times, detailed in table 1. A tornado forms at $t=5580$ seconds, producing a region of low density (figure 3), and a region of high vertical vorticity (seen in figure 4).
@@ II.3 Muon Simulation
MCEq [^1] https://github.com/mceq-project/MCEq/tree/master/MCEq is used to numerically solve the atmospheric particle cascade equations describing the muon flux as the shower propagates through the atmosphere. The primary cosmic ray flux is simulated according to the model described in [21]. Primary cosmic ray interactions are then simulated using the SIBYLL 2.3c interaction model [22].
@@
where $\Delta\theta$ and $\Delta\phi$ correspond to the size of the angular bins in azimuth and elevation (2 degrees). The bounds of the energy integral, $E_{min}$ and $E_{max}$ describe the minimum and maximum muon energies that were simulated. For this paper, muon fluxes were calculated between $E_{min}=0\text{ TeV}$ and $E_{max}=10^{8}\text{ TeV}$, however the vast majority ($>99\%$) of the atmospheric muon flux reaching the ground exists between 0 and 1 TeV.
We choose a total integration time of $\Delta t$ = 1 hour, examining the storm at set times using simulated muon data from the preceding hour. The muon flux is integrated in tandem with the simulated storm evolution, with temporal bin edges defined by the times listed in table 1. The density field is considered to be fixed within each temporal bin, though all bins are smaller than the total integration interval of 1 hour.
To give a sense of scale: for a 1000 square meter muon detector, the expected rate of atmospheric muons is approximately 28,400 events per square degree angular bin per hour at $\theta=40^{\circ}$. This rate varies naturally as a function of elevation angle, with a lower flux near the horizon and a higher flux from very vertical elevation angles [27].
@@ III Results
Figure 2 shows the idealized expectation of the effect of the storm density field on the atmospheric muon flux along various lines of sight for a 1000 m2 muon detector, over the course of 1 hour preceding the time listed for the map. These maps show the observed atmospheric muon flux as a function of azimuthal angle ($\phi$), and elevation angle ($\theta$). $\phi=0^{\circ}$ is defined to point parallel to the positive x-axis, and increases in a counterclockwise direction (so $\phi=90^{\circ}$ is parallel to the positive y-axis). Differences in $\phi$ and $\theta$ can be converted to approximate horizontal and vertical lengths given a distance to the region of interest using trigonometry:
$d_{horizontal}=2R\text{tan}(\Delta\phi)$ (9)
$d_{vertical}=2R\text{tan}(\Delta\theta)$ (10)
where $d_{horizontal}$ and $d_{vertical}$ are the horizontal and vertical distances between two points, $R$ is the distance between the observer and the region of interest, and $\Delta\phi$ and $\Delta\theta$ are differences in $\phi$ and $\theta$ for two arbitrary points on the maps shown in figures like figure 2.
Regions of increased muon flux, identifiable as groups of pixels with lower values of $L$, correspond to low-density regions of the storm. Similarly, regions of suppressed muon flux correspond to higher-density regions. Note, however, that the muon flux is correlated with the average of the density field along a particular line of sight, meaning information about the density field as a function of depth is not easily accessible from the two dimensional images shown. Depth-dependent information about the density field could conceivably be obtained by multiple detectors with different views of the same storm, or by examining the energy distribution of detected muons, however development of this idea is left for future study.
For comparison, plots showing the density perturbation field integrated along lines of sight leading to the muon detector are shown in figure 3. The large region of low density located near $\phi=100$ degrees corresponds to the storm’s mesocyclone. While the mesocyclone is present during all stages of the storm’s lifecycle, the intensification of this region between $t=5580$ and $t=10200$ seconds corresponds to the presence of a tornado. This intensification is reflected in the muon flux, as can be seen in plots of the idealized muon flux in figure 2.
Figure 3: The density perturbation field, relative to the no-storm base state, integrated along line of sight leading to the muon detector (here, $d\ell$ is the line element along the line of sight), at various times during the simulated storm evolution. The tornado forms at $t=5580$ seconds, and is located within the mesocyclone producing the density perturbation near $\phi=100$ degrees.
@@
The plots show in figure 2 show the idealized expectation of the effect of the simulated supercell thunderstorm on the atmospheric muon flux, neglecting statistical variations in the atmospheric muon flux. To improve upon this, Poisson variations are added to $N_{obs}$ and $L$ is recalculated using the new values. Examples of maps of $L$ containing these variations can be seen in figure 5. Near the horizon, the muon flux is lower and the data is dominated by Poisson variations. At high elevations ($\theta>30$ degrees) the structure of the high and low density regions of the storm begin to become visible as clusters of pixels with lower/elevated $L$ values. The intensification of the low density core temporally coincident with tornado formation is still visible, even when accounting for statistical variations in the observed muon flux over a 1 hour time window.
The effect of the tornado can also be seen in the total summed value of the maps shown in figure 5. Low density perturbations will introduce an excess of muon events seen by the detector, resulting in lower per-pixel $\log(L)$ values on average. Figure 6 shows the value of the summed muon images as a function of time during the supercell thunderstorm’s life span. Tornado formation introduces a drop in the spatially integrated muon flux beyond the effect seen from the mesocyclone alone.
@@ IV.1 Muon Detector Size
Smaller detectors may still be able to measure atmospheric density effects on the muon flux. Since variations in the atmospheric muon flux due to supercell thunderstorms are on the order of 1%, muon detectors seeking to measure this effect need to be large enough to collect O($10^{6}$) atmospheric muon counts over the angular area of interest during the lifetime of the storm (typically 10s to 100s of minutes). Measurements taken with smaller detectors will have larger relative statistical fluctuations, making it more difficult to identify the underlying structure of the density field. On the other hand, smaller detectors also have the potential to be significantly more portable. While a 1000 m2 would likely need to be a permanent, stationary installation, a 100 m2 detector could conceivably be moved to a location with projected severe weather several days beforehand, and a 50 m2 detector is small enough to fit on roughly 5 trucks, allowing for the detector to be moved to the location of a severe weather event as it is developing. Examples of muon maps produced by a 100 m2 detector can be seen in figure 7. Though statistical variations conceal much of the structure seen in figure 5, the aggregate effect of the tornado on the underlying density field can still be observed in the sum of the individual pixel values, albeit to a much smaller degree. This is shown in figure 8.
Figure 7: Maps of the muon flux, similar to those shown in figure 5, but as seen by a 100 m2 muon detector. The effect of the low-density core of the tornado is significantly smaller relative to statistical fluctuations in comparison to the 1000 m2 case, though is still faintly visible.
@@ IV.2 Range and Resolution
Though individual density measurements using this technique may not be as precise in measuring the density at a particular location when compared to more traditional in-situ methods, there is a distinct advantage in being able to remotely measure the density field over large distance scales. Measurements of the atmospheric muon flux can reveal unique information of the variation of the density field as a function of height that is inaccessible to ground-based sensors. Additionally, the lack of a need for in-situ instrumentation makes this measurement more logistically straightforward than current methods, as it does not require detectors to be within the direct path of an active tornado.
The range of this technique depends on the desired spatial resolution of the density perturbation field. The atmospheric muon flux naturally varies with elevation angle, with a higher proportion of muon events incident at large elevation angles [27]. For a fixed detector size, a more detailed scan of the density field requires a larger number of muon counts, which can be obtained by using the muon flux a higher (more vertical) elevation angles. However, using data at higher elevation angles reduces the maximum possible horizontal distance between the detector and the storm.
@@
where $d$ is the distance scale over which density variations are measured, $\theta$ is the elevation angle of the target region, and $\Delta\psi$ is obtained for a particular elevation angle from the curves shown in figure 9. Due to the competing effects of the increased resolution at low elevation angles, and the decreased horizontal distance when examining high elevation angles, the horizontal range of this technique peaks near an elevation angle of 45 degrees. This can be seen for various detector sizes and desired distance resolutions in figure 10. Note that here “horizontal range” is the distance along the surface of the earth, accounting for the earth’s curvature.
With a 1000 m2 muon detector, a 5 kilometer resolution measurement of the density field could be obtained at a distance of up to 35 kilometers, though this is restricted to along the $\theta=40^{\circ}$ line of sight. If a larger, permanent muon detector were placed in a tornado-prone location, this would translate to being able to measure the density field of $O(1)$ tornadoes every year [30]. Smaller detectors would have a reduced, but still significant range, capable of measuring density variations over 5 kilometer length scales at a range of 10 kilometers. Such detectors would have the important advantage of portability, with the potential to be deployed or moved near regions expected to experience severe weather in the near future.
@@
Figure 10: Projected horizontal ranges of using this technique to measure density field variations along the listed length scales ($d$). Large elevation angles correspond to a higher muon rate allowing for more precise measurements, but also geometrically correspond to a smaller horizontal distance to a particular altitude layer. These two competing effects are roughly balanced near an elevation angle of 45 degrees, where the horizontal range of this technique is maximized.
@@ IV.4 Other Weather Phenomena
In the more general case, atmospheric phenomena can be characterized by their physical size and the scale of their density perturbations. By simulating various atmospheric phenomena described by these two variables, we can estimate what types of atmospheric phenomena are candidates for remote density measurements using atmospheric muons.
A simplified simulation of the density field is used to achieve this. Different atmospheric phenomena are simulated as 10-kilometer tall cylinders of negative density perturbation, with variable horizontal radius. The radius and average density perturbation are varied, and for each case the muon flux residual integrated over the full field of view is recorded. This can be used to estimate how large of an effect various atmospheric phenomena might have on the muon flux. These curves, as well as highlighted regions corresponding to different atmospheric phenomena, can be seen in figure 12. These results seem consistent with previous studies [10, 12], which have observed perturbations in the muon flux in the vicinity of typhoons (hurricanes) and non-tornadic thunderstorms.
Figure 12: Various severe weather phenomena plotted in the space of size (x-axis), and magnitude of associated density drop (y-axis) [31, 32, 14, 33, 34]. Highlighted regions for tornadoes, thunderstorms, and hurricanes are approximate, as these values have not been exhaustively cataloged for the majority of storms. For the purposes of this plot, supercell thunderstorms are estimated to have density deviations comparable to that of tornadoes, though this has yet to be measured. Curves corresponding to regions of a 1% and 5% effect on the atmospheric muon flux are shown, where points above these curves are expected to produce a larger deviation in the muon flux.